OpenStudy (anonymous):

Which of the following is true? A. The trapezoid and rectangle have the same area and perimeter. B. The trapezoid and rectangle have a different area and perimeter. C. The trapezoid and rectangle have the same area, but different perimeter. D. The trapezoid and rectangle have the same perimeter, but different area.

OpenStudy (anonymous):

The area of trapezoid is 4*(8+2)/2=20 The perimeter is 5+5+8+2=20 The area of rectangle is 4*6=24, perimeter equals 4+4+6+6=20 So the answer is D

OpenStudy (anonymous):

ok thank you so much